OpenStudy (mathlegend):

Solve for q:

OpenStudy (mathlegend):

\[p^{2}q-3q=14\]

OpenStudy (anonymous):

step 1: factor out the "q" on the left

OpenStudy (anonymous):

q ( p^2 - 3) = 14

OpenStudy (mathlegend):

Can't I just combine like terms?

OpenStudy (anonymous):

Then divide left and right by (p^2 - 3)

OpenStudy (anonymous):

you need to get "q" all by itself...

OpenStudy (anonymous):

you could combine them, but you don't know anything about "p" to allow you to combine them more than the problem started.

OpenStudy (mathlegend):

oh

OpenStudy (mathlegend):

Why did you put q(p^2-3) ???

OpenStudy (mathlegend):

Shouldn't it be q(1-3)

OpenStudy (anonymous):

p^{2}q-3q=14 q( p^2 - 3) = 14 <<<--- factor out q q = 14 / (p^2 - 3) <<<<------ divide both sides by (p^2 - 3) to get q alone

OpenStudy (anonymous):

no, q(1-3) would be equal to q-3q you have q(p^2) - 3q

OpenStudy (mathlegend):

oh ok

OpenStudy (mathlegend):

Thank You so much!

OpenStudy (anonymous):

very glad to help :)